### v1

# Patch-wise Structural Loss for Time Series Forecasting

Abstract
Time-series forecasting has gained significant attention in machine learning due to its crucial role in various domains. However, most existing forecasting models rely heavily on point-wise loss functions like Mean Square Error, which treat each time step independently and neglect the structural dependencies inherent in time series data, making it challenging to capture complex temporal patterns accurately. To address these challenges, we propose a novel Patch-wise Structural (PS) loss, designed to enhance structural alignment by comparing time series at the patch level. Through leveraging local statistical properties, such as correlation, variance, and mean, PS loss captures nuanced structural discrepancies overlooked by traditional point-wise losses. Furthermore, it integrates seamlessly with point-wise loss, simultaneously addressing local structural inconsistencies and individual time-step errors. PS loss establishes a novel benchmark for accurately modeling complex time series data and provides a new perspective on time series loss function design. Extensive experiments demonstrate that PS loss significantly improves the performance of state-of-the-art models across diverse real-world datasets. The data and code are publicly available at: https://github.com/Dilfiraa/PS_Loss.
Time Series Forecasting, Loss Functions, Deep Learning
1 Introduction
Time series forecasting plays a crucial role across various real-world domains, including traffic (Kong et al., 2024; Long et al., 2024), weather (Lam et al., 2023; Wu et al., 2023b), and finance (Huang et al., 2024a), where accurate forecasting is essential for informed decision-making. Recent advancements in deep learning have developed models that capture complex dependencies and intricate patterns in time series data, significantly improving forecasting accuracy (Zhou et al., 2021, 2022; Sun et al., 2024a; Yu et al., 2024; Sun et al., 2024b). These models aim to produce multi-step predictions based on historical data that closely follow the trajectory of future series.
(a) Poor correlation
(b) Variance mismatch
(c) Mean offset
Figure 1: Limitations of MSE loss in capturing structural differences. Predictions with similar MSE but different statistical properties: (a) Poor linear correlation leads to misalignment in directionality and pattern; (b) Variance mismatch distorts the fluctuation dynamics; (c) Mean offset causes a shift in the overall level of the forecast.
Despite these advancements, most current forecasting models rely on Mean Squared Error (MSE) as the primary loss function. MSE evaluates predictions by averaging point-wise distances, which overlooks important structural characteristics inherent in time series data, leading to suboptimal forecasts that fail to accurately capture the underlying temporal dynamics (Le Guen & Thome, 2019; Lee et al., 2022; Cuturi & Blondel, 2017).
To illustrate the limitation of MSE, we consider three predictions with similar MSE values but different forecasting quality, as shown in Figure 1. The prediction in Figure 1(a) exhibits poor correlation with the ground truth, failing to capture the overall directionality and pattern. In contrast, the prediction in Figure 1(b) aligns with the general shape of the ground truth but fails to match its variability, as indicated by the variance mismatch, leading to distorted fluctuation dynamics. On the other hand, the prediction in Figure 1(c) maintains the shape but exhibits a mean offset, causing a consistent shift in the overall forecast level. These examples highlight that correlation, variance, and mean capture distinct yet complementary aspects of structural similarity, providing a more comprehensive assessment of the alignment between prediction and ground truth.
However, measuring correlation, mean, and variance at a global level is insufficient, as global metrics often fail to capture local variations in statistical properties, especially in non-stationary time series where these properties evolve over time. This limitation highlights the need for a localized approach that can adapt to these variations and provide a more precise comparison of time series.
To address these challenges, we propose a Patch-wise Structural (PS) loss that improves time series forecasting by introducing structural similarity metrics at the patch level. Our method comprises three key components. First, we introduce the Fourier-based Adaptive Patching (FAP) module, which adaptively segments the time series into patches, enabling a finer-grained assessment of structural alignment. Second, we propose the Patch-wise Structural (PS) loss, which integrates correlation, variance, and mean losses to capture local structural similarities and account for variations across different regions of the data. Third, we employ a Gradient-based Dynamic Weighting (GDW) strategy to adaptively balance the contributions of each loss term based on their gradients, ensuring that all structural aspects are adequately accounted for during training. PS loss can be seamlessly combined with MSE loss, resulting in more robust and accurate forecasting. Our contributions are as follows:
  • We introduce a novel structural similarity measure that combines correlation, variance, and mean, offering a more comprehensive understanding of structural patterns and improving alignment between predictions and ground truth.
  • We propose a patch-wise comparison method that enables localized similarity measures and accounts for changing statistical properties within the series, ensuring accurate structural alignment.
  • Extensive experiments on five state-of-the-art architectures and seven real-world datasets confirm the effectiveness of our approach, demonstrating consistent performance gains and enhanced forecasting accuracy.
2 Related Work
2.1 Time Series Forecasting Models
In recent years, numerous deep forecasting models have been proposed to capture complex dependencies in time series data by leveraging the powerful representation capabilities of neural networks (Kim et al., 2024). Based on their architecture, these models can be categorized into CNN, Transformer, MLP, and LLM-based approaches. CNN-based models excel at modeling local temporal relationships (Wu et al., 2023a; Luo & Wang, 2024; Cheng et al., 2024; Li et al., 2024; Liu et al., 2022). In contrast, Transformer-based models have gained significant attention for their ability to model long-range dependencies through self-attention mechanisms (Vaswani et al., 2017; Nie et al., 2023; Liu et al., 2024a; Chen et al., 2024b; Lin et al., 2024b; Zhou et al., 2022, 2021). Additionally, MLP-based models offer efficient alternatives that balance simplicity with performance (Zeng et al., 2023; Wang et al., 2024b; Yi et al., 2024; Lin et al., 2024a; Ekambaram et al., 2023). Recently, LLM-based models have emerged as a promising direction, leveraging the capabilities of large language models to effectively capture complex temporal dependencies (Zhou et al., 2023; Liu et al., 2024b). These models use point-wise MSE as their loss function, treating each time step independently, which limits their ability to capture the structural nuances of time series data and may hinder forecasting performance.
Figure 2: Overview of the proposed PS loss. The method consists of three main components: (1) Fourier-based Adaptive Patching, where the ground truth $Y$ and predicted series $\hat{Y}$ are adaptively segmented into patches; (2) Patch-wise Structural Loss, which measures local similarity between patches by integrating correlation ($\mathcal{L}_{Corr}$), variance ($\mathcal{L}_{Var}$), and mean losses ($\mathcal{L}_{Mean}$); and (3) Gradient-based Dynamic Weighting, which dynamically adjusts the weights of these loss components ($\alpha$, $\beta$, and $\gamma$) based on their gradient magnitudes to ensure balanced optimization. The PS loss ($\mathcal{L}_{PS}$) is seamlessly integrated to the MSE loss ($\mathcal{L}_{MSE}$) to improve forecasting accuracy.
2.2 Loss Functions for Time Series Forecasting
Recent efforts to address the limitations of MSE have given rise to alternative loss functions (Le Guen & Thome, 2019; Cuturi & Blondel, 2017; Xiong et al., 2024; Qi et al., 2024), which can be broadly categorized into shape-focused losses and dependency-focused losses. Shape-focused losses aim to capture structural similarities between the ground truth and predictions by explicitly addressing shape misalignment. For instance, Dynamic Time Warping (DTW)-based methods such as Soft-DTW (Cuturi & Blondel, 2017) and DILATE (Le Guen & Thome, 2019) align sequences under temporal distortions. While effective at improving shape alignment, their high computational complexity limits scalability. TILDE-Q (Lee et al., 2022), on the other hand, introduces transformation invariance, making it robust to amplitude shifts, phase shifts, and scaling differences, thus focusing on shape similarity. Dependency-focused losses focus on capturing temporal dependencies within the forecast window. FreDF (Wang et al., 2024a) bypasses label correlation complexity by learning to forecast in the frequency domain. Despite these advancements, most prior methods rely on global comparisons of entire time series, overlooking crucial localized structural details. In contrast, our proposed approach incorporates patch-wise statistical properties into the loss function, enabling a more granular structural measurement of the data and providing a fundamentally new perspective on time series loss design.
3 Methodology
3.1 Preliminaries
Given a historical time series $\mathbf{X}=\{x_{0},x_{1},\ldots,x_{L-1}\}\in\mathbb{R}^{C\times L}$, where $C$ is the number of channels and $L$ is the length of the lookback window. The goal of time series forecasting is to learn a mapping $g:\mathbb{R}^{C\times L}\rightarrow\mathbb{R}^{C\times T}$ that generates predictions $\hat{\mathbf{Y}}=\{\hat{y}_{0},\hat{y}_{1},\ldots,\hat{y}_{T-1}\}\in\mathbb{R}^{C\times T}$ approximating the ground truth $\mathbf{Y}=\{y_{0},y_{1},\ldots,y_{T-1}\}\in\mathbb{R}^{C\times T}$, where $T$ is the forecasting length. For simplicity of notation, we focus on the univariate case ($C=1$) in the following discussions. MSE is one of the most common loss functions for evaluating prediction accuracy, and it is defined as:
$\mathcal{L}_{MSE}=\frac{1}{T}\sum_{i=0}^{T-1}(y_{i}-\hat{y}_{i})^{2}.$ (1)
3.2 Overview
Our method addresses the structural alignment limitations of MSE loss by focusing on localized pattern differences. To achieve this, we incorporate three key components: Fourier-based Adaptive Patching (FAP), Patch-wise Structural (PS) Loss, and Gradient-Based Dynamic Weighting (GDW), which collectively measure the patch-wise structural similarity between prediction and ground truth, as illustrated in Figure 2.
We first generate predictions $\hat{\mathbf{Y}}$ using a backbone model, forming the basis for subsequent patch-level comparisons. Following this, FAP segments both the ground truth and prediction into patches based on the dominant frequency of the ground truth, ensuring that the patches capture recurring patterns and coherent structural information. Once the series is segmented, the PS loss evaluates the similarity between patches by integrating correlation, variance, and mean losses, enabling a comprehensive assessment of local structural alignment across the time series. To further refine the optimization process, GDW dynamically adjusts the weights of the three losses based on their gradients, ensuring balanced contributions from each component and leading to more effective optimization and consistent model training.
By seamlessly combining PS loss with MSE, we handle both global point-wise errors and local structural discrepancies. This integration enhances forecasting accuracy, demonstrating the advantages of incorporating patch-wise structural analysis into time series loss functions.
3.3 Fourier-based Adaptive Patching
To enable localized structural similarity measures between time series, we propose a patch-wise comparison approach that divides the prediction $\hat{\mathbf{Y}}\in\mathbb{R}^{T}$ and the ground truth $\mathbf{Y}\in\mathbb{R}^{T}$ into patches. By comparing these patches, we can better account for the structural patterns within the series.
In order to adaptively determine the patch length, we introduce the Fourier-based Adaptive Patching (FPA) technique. By analyzing the amplitude spectrum of the ground truth via the Fast Fourier Transform (FFT), we identify its dominant periodicity:
$\mathbf{A}=\operatorname{Amp}\left(\mathcal{FFT}(\mathbf{Y})\right),\quad f=\underset{f_{*}\in\{1,\ldots,\lceil T/2\rceil\}}{\arg\max}(\mathbf{A}),$ (2)
where, $\operatorname{Amp}(\cdot)$ calculates the amplitude values and $f$ denotes the dominant frequency. The period $p=\lceil T/f\rceil$ serves as an initial patch length candidate, ensuring the patches reflect recurring structural patterns. However, dominant periodicity alone may produce overly large patches, missing finer structure nuances. To balance granularity and coherence, the final patch length is determined as:
$P=\min\left(\left\lceil\frac{p}{2}\right\rceil,\delta\right),\quad p=\left\lceil\frac{T}{f}\right\rceil,$ (3)
where $\delta$ is a predefined threshold. This adjustment ensures patches remain focused on localized structural patterns, even when the dominant periodicity is unclear or too large.
With patch length $P$ and stride $S=\lceil P/2\rceil$, the series is segmented into patches $\hat{\mathbf{Y}}_{p}\in\mathbb{R}^{P\times N}$ and $\mathbf{Y}_{p}\in\mathbb{R}^{P\times N}$, where $N=\lfloor(T-P)/S\rfloor+1$ is the total number of patches. The $i$-th patch of $\mathbf{Y}$ is defined as $\mathbf{Y}_{p}^{(i)}=\{y_{0}^{(i)},y_{1}^{(i)},\ldots,y_{P-1}^{(i)}\}$, where $y_{j}^{(i)}$ refers to the $y_{i\times S+j}$ in the unpatched time series.
3.4 Patch-wise Structural Loss
Building on the adaptive patching framework, we propose the Patch-wise Structural (PS) loss to enhance time series forecasting by capturing and aligning localized statistical properties within each patch. PS loss focuses on three key metrics: correlation, variance, and mean - to offer a deeper structural comparison.
Correlation Loss. Correlation quantifies the degree of directional and pattern consistency between the ground truth and predicted patches. We define the correlation loss based on the Pearson Correlation Coefficient (PCC) (Cohen et al., 2009), as follows:
$\mathcal{L}_{Corr}$ $=\frac{1}{N}\sum_{i=0}^{N-1}1-\rho(Y_{p}^{(i)},\hat{Y}_{p}^{(i)})$ (4)
$=\frac{1}{N}\sum_{i=0}^{N-1}1-\frac{\sum_{j=0}^{P-1}(y_{j}^{(i)}-\mu^{(i)})(\hat{y}_{j}^{(i)}-\hat{\mu}^{(i)})}{\sigma^{(i)}\hat{\sigma}^{(i)}},$
where $\mathbf{\mu}^{(i)}$ and $\mathbf{\sigma}^{(i)}$ represent the mean and standard deviation of the $i$-th patch, respectively, and $\rho(\cdot,\cdot)$ denotes the PCC. Minimizing $\mathcal{L}_{Corr}$ encourages the model to align its predictions with the ground truth’s directionality and pattern trends. This alignment helps the model capture underlying pattern relationships, which are critical for time series forecasting tasks that require more than just point-wise accuracy.
Variance Loss. Variance quantifies the degree of fluctuation within a patch, providing a localized measure of variability. To ensure that predictions represent these localized dynamics, we introduce a variance loss that compares the relative dispersion between ground truth and predicted patches.
Rather than directly comparing variance values, our approach focuses on softly aligning the relative dispersion of each time step within patches. Specifically, we compute the deviations, $Y_{p}^{(i)}-\mu^{(i)}$ and $\hat{Y}_{p}^{(i)}-\hat{\mu}^{(i)}$, and normalize them into probability distributions using the softmax function. This transformation highlights the relative magnitude of the deviations, allowing us to focus on how the values within each patch are dispersed relative to one another (Shu et al., 2021). The similarity between these distributions is measured via the Kullback–Leibler (KL) divergence, leading to the following variance loss:
$\mathcal{L}_{Var}=\frac{1}{N}\sum_{i=0}^{N-1}\operatorname{KL}\left(\phi\left(Y_{p}^{(i)}-\mu^{(i)}\right)\,\|\,\phi\left(\hat{Y}_{p}^{(i)}-\hat{\mu}^{(i)}\right)\right),$ (5)
where $\phi(\cdot)$ is the softmax function. Due to the shift-invariance property of softmax, i.e., $\phi(Y_{p}^{(i)}-\mu^{(i)})=\phi(Y_{p}^{(i)})$, the variance loss formulation simplifies to:
$\mathcal{L}_{Var}=\frac{1}{N}\sum_{i=0}^{N-1}\operatorname{KL}\left(\phi\left(Y_{p}^{(i)}\right)\,\|\,\phi\left(\hat{Y}_{p}^{(i)}\right)\right).$ (6)
By aligning the relative dispersion within patches, the variance loss encourages the model to capture the underlying fluctuation patterns of the time series, leading to more structurally coherent forecasts.
Mean Loss. The mean value of a patch provides a clear indicator of its central tendency. While correlation and variance losses effectively capture directionality patterns and localized variability, they may overlook systematic shifts or biases in the predictions. To address this, we introduce a mean loss term that directly measures the consistency of average values.
Specifically, the mean loss is defined as the Mean Absolute Error (MAE) between the mean values of the ground truth and predicted patches:
$\mathcal{L}_{Mean}=\frac{1}{N}\sum_{i=0}^{N-1}\left|\mu^{(i)}-\hat{\mu}^{(i)}\right|,$ (7)
where $\mu^{(i)}$ and $\hat{\mu}^{(i)}$ represent the mean values of the $i$-th patch in the ground truth and predictions, respectively. By aligning the mean values, the mean loss corrects prediction biases and ensures the model maintains accurate overall levels and central tendencies, enabling it to adapt to dynamic shifts in the time series.
PS Loss. Finally, we integrate the correlation, variance, and mean losses into a unified Patch-wise Structural (PS) loss, which is formulated as the weighted sum of the three individual loss components:
$\mathcal{L}_{PS}=\alpha\mathcal{L}_{Corr}+\beta\mathcal{L}_{Var}+\gamma\mathcal{L}_{Mean},$ (8)
where $\alpha$, $\beta$, and $\gamma$ balance the contributions of the individual components. This integration provides a comprehensive structural alignment by leveraging the strengths of each component loss.
3.5 Gradient-based Dynamic Weighting
PS loss combines three loss components to measure patch-level similarity. To ensure effective training, it is crucial to balance these components, preventing any single loss term from dominating the optimization process (Chen et al., 2018). In multi-task learning, a common approach to balancing multiple losses is to weight them based on their gradient magnitudes, which is robust to differences in loss scales and directly reflects each loss’s influence on the optimization process (Sener & Koltun, 2018; Lin et al., 2019; Liu et al., ). Inspired by this, we propose a Gradient-based Dynamic Weighting (GDW) strategy, which adaptively adjusts loss weights based on gradients to achieve balanced optimization.
At training step $t$, let $W$ denote the model’s output layer parameters. We compute the L2-norm of each loss component’s gradient with respect to $W$:
$G_{Corr}^{(t)}$ $=\|\nabla_{W}\mathcal{L}_{Corr}^{(t)}\|_{2},\quad G_{Var}^{(t)}=\|\nabla_{W}\mathcal{L}_{Var}^{(t)}\|_{2},$ (9)
$G_{Mean}^{(t)}$ $=\|\nabla_{W}\mathcal{L}_{Mean}^{(t)}\|_{2},$
where $G_{Corr}^{(t)}$ is the gradient magnitude of the correlation loss, and $G_{Var}^{(t)}$, $G_{Mean}^{(t)}$ are defined analogously. Using these gradient magnitudes, we compute the average gradient magnitude:
$\overline{G}^{(t)}=\frac{G_{Corr}^{(t)}+G_{Var}^{(t)}+G_{Mean}^{(t)}}{3}.$ (10)
The loss weights at iteration $t$ are then set as:
$\alpha^{(t)}=\frac{\overline{G}^{(t)}}{G_{Corr}^{(t)}},\quad\beta^{(t)}=\frac{\overline{G}^{(t)}}{G_{Var}^{(t)}},\quad\gamma^{(t)}=c\cdot v\cdot\frac{\overline{G}^{(t)}}{G_{Mean}^{(t)}}.$ (11)
To further refine the mean loss contribution, we introduce two scaling factors, $c$ and $v$, which adjust the weight of the mean loss based on the model’s performance in correlation and variance. These factors are defined as:
$c=\frac{1}{2}\cdot(1+\frac{\sigma_{Y\hat{Y}}}{\sigma_{Y}\sigma_{\hat{Y}}}),\quad v=\frac{2\sigma_{Y}\sigma_{\hat{Y}}}{\sigma_{Y}^{2}+\sigma_{\hat{Y}}^{2}},$ (12)
where $\sigma_{Y\hat{Y}}$ is the covariance between ground truth and predictions, and $\sigma_{Y}$ and $\sigma_{\hat{Y}}$ are their standard deviations. Both $c$ and $v$ lie in the range $[0,1]$, increasing as the predictions align more closely with the ground truth in terms of correlation and variance. As alignment improves, the relative weight of the mean loss increases, ensuring it plays a greater role when correlation and variance are well-matched, thereby enhancing the refinement of predictions.
Finally, the total loss function for optimizing the time series forecasting model is formulated as the sum of the standard MSE loss and the PS loss:
$\mathcal{L}=\mathcal{L}_{MSE}+\lambda\mathcal{L}_{PS},$ (13)
where $\lambda$ is a hyperparameter that controls the contribution of PS loss.
4 Experiments
4.1 Experimental Setup
Table 1: Long-term multivariate forecasting results. The table reports MSE and MAE for different forecasting lengths $T\in\{96,192,336,720\}$. The input sequence length is consistent with the backbone setting. The better results are highlighted in bold.
Models iTransformer (2024a) PatchTST (2023) TimeMixer (2024b) DLinear (2023) TimseNet (2023a)
Loss Functions MSE + PS MSE + PS MSE + PS MSE + PS MSE + PS
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.387 0.405 0.379 0.396 0.375 0.400 0.368 0.394 0.385 0.402 0.366 0.392 0.384 0.405 0.367 0.389 0.384 0.405 0.382 0.406
192 0.441 0.436 0.428 0.424 0.414 0.421 0.407 0.415 0.443 0.430 0.421 0.421 0.443 0.450 0.402 0.411 0.443 0.450 0.424 0.432
336 0.491 0.462 0.474 0.453 0.432 0.436 0.420 0.426 0.512 0.470 0.489 0.453 0.447 0.448 0.435 0.435 0.494 0.471 0.457 0.448
720 0.509 0.494 0.480 0.478 0.450 0.466 0.425 0.449 0.497 0.476 0.474 0.464 0.504 0.515 0.463 0.484 0.504 0.515 0.477 0.477
Avg 0.457 0.449 0.440 0.438 0.418 0.431 0.405 0.421 0.460 0.444 0.437 0.432 0.445 0.454 0.417 0.430 0.457 0.460 0.435 0.441
ETTh2 96 0.301 0.350 0.289 0.340 0.274 0.336 0.272 0.332 0.291 0.342 0.284 0.330 0.290 0.353 0.280 0.341 0.330 0.367 0.316 0.354
192 0.380 0.399 0.373 0.391 0.339 0.380 0.336 0.375 0.376 0.396 0.355 0.381 0.388 0.422 0.358 0.396 0.405 0.415 0.387 0.396
336 0.424 0.432 0.416 0.428 0.331 0.380 0.324 0.377 0.437 0.439 0.412 0.423 0.463 0.473 0.435 0.450 0.454 0.451 0.428 0.431
720 0.430 0.447 0.421 0.441 0.378 0.421 0.376 0.417 0.464 0.464 0.423 0.440 0.733 0.606 0.597 0.540 0.434 0.448 0.438 0.448
Avg 0.384 0.407 0.375 0.400 0.331 0.379 0.327 0.375 0.392 0.410 0.369 0.394 0.469 0.463 0.417 0.432 0.406 0.420 0.392 0.407
ETTm1 96 0.342 0.377 0.326 0.360 0.288 0.342 0.284 0.334 0.328 0.364 0.316 0.350 0.301 0.345 0.296 0.339 0.334 0.375 0.329 0.369
192 0.383 0.396 0.374 0.384 0.334 0.372 0.328 0.364 0.364 0.382 0.357 0.373 0.336 0.366 0.333 0.362 0.406 0.413 0.388 0.400
336 0.418 0.418 0.410 0.406 0.367 0.393 0.357 0.383 0.387 0.402 0.385 0.394 0.372 0.389 0.365 0.380 0.415 0.422 0.394 0.406
720 0.487 0.457 0.472 0.440 0.417 0.422 0.406 0.412 0.472 0.449 0.444 0.432 0.427 0.423 0.419 0.413 0.511 0.472 0.469 0.444
Avg 0.408 0.412 0.396 0.397 0.352 0.382 0.344 0.373 0.388 0.399 0.375 0.387 0.359 0.381 0.353 0.374 0.417 0.421 0.395 0.405
ETTm2 96 0.186 0.272 0.175 0.254 0.164 0.253 0.160 0.247 0.175 0.257 0.171 0.252 0.172 0.267 0.163 0.251 0.189 0.266 0.180 0.257
192 0.254 0.314 0.242 0.299 0.221 0.291 0.216 0.287 0.240 0.302 0.234 0.295 0.237 0.314 0.222 0.296 0.263 0.312 0.246 0.298
336 0.316 0.351 0.304 0.338 0.277 0.329 0.267 0.321 0.303 0.343 0.291 0.331 0.295 0.359 0.277 0.332 0.326 0.354 0.303 0.334
720 0.414 0.407 0.401 0.394 0.365 0.384 0.357 0.378 0.396 0.400 0.386 0.389 0.427 0.439 0.377 0.397 0.418 0.405 0.418 0.402
Avg 0.292 0.336 0.281 0.321 0.257 0.314 0.250 0.308 0.278 0.326 0.270 0.317 0.283 0.345 0.260 0.319 0.299 0.334 0.286 0.323
Weather 96 0.176 0.216 0.167 0.203 0.151 0.198 0.149 0.190 0.165 0.212 0.161 0.201 0.174 0.233 0.171 0.222 0.172 0.220 0.170 0.219
192 0.227 0.260 0.219 0.249 0.196 0.242 0.193 0.235 0.211 0.254 0.207 0.244 0.218 0.278 0.212 0.259 0.225 0.264 0.225 0.263
336 0.282 0.299 0.274 0.291 0.248 0.282 0.245 0.276 0.263 0.293 0.263 0.285 0.263 0.314 0.258 0.300 0.281 0.304 0.281 0.303
720 0.357 0.348 0.353 0.343 0.319 0.335 0.318 0.330 0.343 0.345 0.340 0.337 0.332 0.374 0.323 0.356 0.359 0.354 0.358 0.352
Avg 0.261 0.281 0.253 0.272 0.228 0.264 0.226 0.258 0.245 0.276 0.243 0.267 0.247 0.300 0.241 0.284 0.259 0.285 0.259 0.284
ECL 96 0.148 0.239 0.146 0.236 0.130 0.223 0.128 0.220 0.153 0.245 0.152 0.245 0.140 0.237 0.140 0.236 0.168 0.272 0.164 0.266
192 0.167 0.258 0.161 0.250 0.149 0.240 0.145 0.235 0.166 0.257 0.165 0.256 0.154 0.250 0.153 0.248 0.186 0.289 0.177 0.277
336 0.178 0.271 0.174 0.265 0.165 0.257 0.161 0.252 0.185 0.275 0.182 0.272 0.169 0.268 0.169 0.266 0.196 0.297 0.192 0.292
720 0.211 0.300 0.208 0.297 0.208 0.296 0.196 0.282 0.224 0.312 0.223 0.312 0.204 0.300 0.203 0.297 0.235 0.329 0.220 0.314
Avg 0.176 0.267 0.172 0.262 0.163 0.254 0.158 0.247 0.182 0.272 0.180 0.271 0.167 0.264 0.166 0.262 0.196 0.297 0.188 0.287
Exchange 96 0.086 0.206 0.087 0.206 0.093 0.214 0.089 0.208 0.086 0.204 0.084 0.202 0.085 0.209 0.078 0.199 0.105 0.235 0.107 0.236
192 0.181 0.304 0.180 0.303 0.194 0.315 0.186 0.307 0.187 0.306 0.186 0.306 0.162 0.296 0.156 0.290 0.232 0.351 0.215 0.337
336 0.338 0.422 0.336 0.420 0.355 0.436 0.345 0.427 0.386 0.454 0.332 0.417 0.333 0.441 0.325 0.432 0.393 0.462 0.377 0.450
720 0.853 0.696 0.860 0.700 0.903 0.712 0.904 0.713 0.928 0.727 0.917 0.719 0.898 0.725 0.811 0.690 1.011 0.768 1.010 0.768
Avg 0.364 0.407 0.366 0.407 0.386 0.419 0.381 0.414 0.397 0.423 0.380 0.411 0.381 0.420 0.373 0.412 0.435 0.454 0.427 0.448
Datasets. We conduct our experiments on seven real-world multivariate time series datasets, including ETT (ETTh1, ETTh2, ETTm1, ETTm2), Weather, ECL, and Exchange. For data preprocessing, we follow the standard protocol used by backbone studies. The detailed experimental setup is described in Appendix A.
Backbones. We select five state-of-the-art (SOTA) time series forecasting models with diverse architectures to comprehensively evaluate PS loss function. Specifically, we include Transformer-based models: iTransformer (Liu et al., 2024a), PatchTST (Nie et al., 2023); MLP-based models: DLinear (Zeng et al., 2023), TimeMixer (Wang et al., 2024b); CNN-based model: TimesNet (Wu et al., 2023a).
Implementation Details. We use the official implementations of each backbone from their respective repositories for our experiments. For fair evaluation, when integrating PS loss to enhance the backbone model’s performance, we follow their original experimental and hyperparameter settings, while only tuning the PS loss weight $\lambda$ and the patch length threshold $\delta$. All the experiments are implemented by PyTorch on a single NVIDIA RTX 3090 24GB GPU.
4.2 Main Results
We present the MSE and MAE on seven real-world datasets for five SOTA long-term multivariate forecasting models in Table 1. Notably, PS loss achieves a remarkable improvement on DLinear, with average reductions of 5.22% in MSE and 4.39% in MAE. Similar performance gains are observed across other backbone models and PS loss consistently enhances the forecasting performance significantly outperforming MSE loss in 134 out of 140 cases, further validating the robustness and broad applicability of the proposed loss function. These improvements highlight the unique strength of PS loss in capturing localized structural patterns within time series data.
Table 2: LLM-based long-term multivariate forecasting results. The table reports MSE and MAE for different forecasting lengths $T\in\{96,192,336,720\}$. The input sequence length is consistent with the backbone setting, $L=336$ for OFA and $L=672$ for AutoTimes. The better results are highlighted in bold.
Models Loss Dataset ETTh1 ETTh2 ETTm1 ETTm2
Functions Pred Len 96 192 336 720 96 192 336 720 96 192 336 720 96 192 336 720
MSE MSE 0.382 0.424 0.444 0.445 0.292 0.359 0.379 0.418 0.292 0.359 0.379 0.418 0.298 0.336 0.368 0.419
OFA MAE 0.402 0.425 0.437 0.458 0.348 0.398 0.414 0.454 0.348 0.398 0.414 0.454 0.344 0.375 0.392 0.425
(2023) + PS MSE 0.371 0.413 0.434 0.422 0.283 0.355 0.372 0.398 0.283 0.355 0.372 0.398 0.290 0.331 0.364 0.416
MAE 0.393 0.416 0.430 0.445 0.338 0.381 0.402 0.431 0.338 0.381 0.402 0.431 0.343 0.367 0.388 0.418
MSE MSE 0.360 0.388 0.401 0.405 0.295 0.359 0.392 0.465 0.289 0.337 0.373 0.430 0.181 0.244 0.299 0.372
AutoTmies MAE 0.400 0.419 0.429 0.440 0.355 0.400 0.431 0.483 0.345 0.374 0.396 0.432 0.269 0.311 0.349 0.398
(2024b) + PS MSE 0.358 0.386 0.398 0.404 0.287 0.348 0.370 0.422 0.280 0.327 0.360 0.416 0.171 0.231 0.284 0.366
MAE 0.398 0.415 0.425 0.438 0.347 0.387 0.410 0.450 0.337 0.366 0.387 0.421 0.258 0.299 0.335 0.387
4.3 Results on LLM-based Models
To further evaluate the generalizability and effectiveness of the proposed PS loss, we conduct experiments on two LLM-based time series forecasting models: OFA (Zhou et al., 2023), built on GPT-2 (Radford et al., ) and Autotimes (Liu et al., 2024b), which utilizes LLaMA (Touvron et al., 2023).
The results in Table 2 confirm that incorporating PS loss produces consistent performance improvements over the standard MSE loss. These findings highlight the potential of PS loss to serve as a universal enhancement for LLM-based time series models, further solidifying its value as a powerful tool for advancing multivariate time series forecasting.
Table 3: Comparison between the proposed Patch-wise Structural loss and other loss functions. The model is iTransformer and we report the result of three datasets-ETTh1, ETTm1, and Weather. The best results are highlighted in bold, and the second-best results are highlighted in underline.
Dataset ETTh1 ETTm1 Weather
Forecast length 96 192 336 720 96 192 336 720 96 192 336 720
MSE MSE 0.387 0.441 0.491 0.509 0.342 0.383 0.418 0.487 0.176 0.227 0.282 0.357
MAE 0.405 0.436 0.462 0.494 0.377 0.396 0.418 0.457 0.216 0.260 0.299 0.348
TILDE-Q MSE 0.386 0.434 0.477 0.521 0.336 0.382 0.418 0.483 0.172 0.221 0.279 0.356
(2022) MAE 0.401 0.429 0.454 0.502 0.366 0.392 0.416 0.453 0.208 0.252 0.296 0.346
FreDF MSE 0.381 0.433 0.471 0.487 0.334 0.380 0.416 0.486 0.170 0.222 0.279 0.357
(2024a) MAE 0.396 0.426 0.446 0.481 0.367 0.393 0.414 0.451 0.210 0.254 0.297 0.348
+ PS MSE 0.379 0.428 0.474 0.480 0.326 0.374 0.410 0.472 0.167 0.219 0.274 0.353
MAE 0.396 0.424 0.453 0.478 0.360 0.384 0.406 0.440 0.203 0.249 0.291 0.343
Table 4: Ablation study of the components of PS loss on the ETTh1 and Weather datasets using DLinear as a backbone.
Method Dlinear + PS w/o $\mathcal{L}_{Corr}$ w/o $\mathcal{L}_{Var}$ w/o $\mathcal{L}_{Mean}$ w/o Patching w/o Weighting
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.367 0.389 0.381 0.402 0.380 0.399 0.372 0.395 0.380 0.401 0.384 0.404
192 0.402 0.411 0.439 0.446 0.440 0.444 0.404 0.414 0.436 0.441 0.403 0.411
336 0.435 0.435 0.443 0.445 0.439 0.486 0.440 0.443 0.442 0.442 0.439 0.440
720 0.463 0.484 0.510 0.519 0.509 0.517 0.474 0.493 0.522 0.528 0.528 0.532
Weather 96 0.171 0.222 0.174 0.233 0.172 0.223 0.172 0.223 0.171 0.224 0.172 0.226
192 0.212 0.259 0.222 0.286 0.217 0.271 0.217 0.271 0.220 0.280 0.214 0.268
336 0.258 0.300 0.261 0.311 0.259 0.303 0.259 0.303 0.263 0.314 0.261 0.307
720 0.323 0.356 0.329 0.369 0.328 0.365 0.328 0.365 0.331 0.372 0.332 0.370
4.4 Comparison with Other Loss Functions
We evaluate the proposed PS loss against advanced loss functions. TILDE-Q (Lee et al., 2022) emphasizes shape similarity using transformation-invariant loss terms, while FreDF (Wang et al., 2024a) bypasses label correlation by comparing time series in the frequency domain. The results are shown in Table 3. PS loss achieves the lowest MSE and MAE in most cases across various datasets and forecasting horizons. This is due to its ability to effectively measure localized structural similarity by evaluating correlation, variance, and mean at the patch level, which enables a more precise alignment between the ground truth and predictions.
4.5 Ablation Study
To evaluate the contribution of each component within the proposed PS loss, we conducted ablation experiments on the ETTh1 and weather datasets using DLinear and iTransformer as the backbone models. The results are presented in Table 4, with additional results provided in Appendix H.
Loss components. Removing any one of the three loss components leads to noticeable performance degradation across all horizons. This highlights the synergistic relationship between these components in capturing localized structural similarities.
Adaptive patching. To evaluate the importance of patching, we conducted a direct comparison by processing the entire time series without segmentation. This approach significantly worsens performance, especially for longer horizons, underscoring the value of the patching mechanism in capturing fine-grained patterns.
Dynamic weighting. To evaluate the importance of dynamic weighting, we replaced the adaptive weighting scheme with a fixed weight of 1 for all three loss terms. This static weighting strategy results in performance degradation, demonstrating that dynamic weighting effectively balances each loss term and improves optimization and forecasting.
Table 5: Zero-shot forecasting results on ETT datasets. The table reports MSE and MAE values for backbone models trained with and without PS loss for zero-shot forecasting. The forecasting length is 192.
Models iTransformer DLinear TimeMixer
Loss Function MSE loss + PS loss MSE loss + PS loss MSE loss + PS loss
Source Target MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 ETTh2 0.379 0.395 0.381 0.394 0.464 0.465 0.356 0.394 0.378 0.391 0.378 0.390
ETTm1 0.963 0.611 0.798 0.575 0.737 0.571 0.717 0.550 1.015 0.633 0.923 0.606
ETTm2 0.299 0.353 0.280 0.338 0.383 0.436 0.273 0.353 0.290 0.348 0.285 0.343
ETTh2 ETTh1 0.617 0.538 0.541 0.498 0.446 0.446 0.436 0.437 0.667 0.552 0.645 0.541
ETTm1 0.924 0.607 0.915 0.599 0.713 0.551 0.732 0.559 1.081 0.635 0.991 0.619
ETTm2 0.292 0.350 0.291 0.349 0.307 0.388 0.268 0.346 0.296 0.353 0.296 0.352
ETTm1 ETTh1 0.718 0.564 0.601 0.516 0.479 0.459 0.472 0.453 0.774 0.576 0.628 0.524
ETTh2 0.443 0.436 0.435 0.427 0.375 0.411 0.371 0.407 0.492 0.459 0.466 0.443
ETTm2 0.260 0.314 0.260 0.310 0.244 0.322 0.239 0.315 0.262 0.315 0.257 0.307
ETTm2 ETTh1 0.914 0.637 0.613 0.523 0.542 0.498 0.498 0.470 0.622 0.530 0.590 0.514
ETTh2 0.447 0.445 0.426 0.421 0.385 0.410 0.366 0.397 0.415 0.424 0.412 0.417
ETTm1 0.674 0.525 0.562 0.476 0.417 0.420 0.375 0.394 0.659 0.514 0.523 0.460
4.6 Zero-shot Forecasting
To evaluate how PS loss improves generalization to unseen datasets, we conducted zero-shot forecasting experiments. Following prior works (Jin et al., 2024; Chen et al., 2024a), we sequentially use each of ETTh1, ETTh2, ETTm1, and ETTm2 as the source dataset, treating the remaining datasets as targets.
The results in Table 5, measured on target datasets with a forecasting length of 192, highlight the consistent advantages of PS loss. It outperforms MSE loss in 34 out of 36 scenarios, demonstrating substantial improvements in generalization across diverse datasets and granularities. These gains arise from PS loss’s ability to analyze time series at the patch level, incorporating local statistical properties to achieve better structural alignment and enabling models to adapt more effectively to unseen data distributions.
(a) Better Correlation
(b) Better Variance Match
(c) Better Mean Match
Figure 3: Forecasting visualization comparing PS loss and MSE loss as objective functions. Predictions with PS loss align more closely with the ground truth’s statistical properties, yielding results with higher structural similarity.
4.7 Forecasting Visualization
We visualize samples from the Weather dataset using iTransformer as the backbone model. As shown in Figure 3, PS loss significantly improves alignment with the ground truth compared to MSE loss: (a) its predictions closely follow the trajectory of the ground truth, preserving overall trends and patterns. (b) it effectively captures amplitude fluctuations, reflecting the localized variability. (c) it reduces biases at the overall level, ensuring better alignment with the central tendency of the ground truth. These results highlight the strengths of PS loss in capturing localized structural patterns. More visualizations are provided in Appendix G.
4.8 Hyperparameter Sensitivity
Our method introduces two key hyperparameters: PS loss weight $\lambda$ and patch length threshold $\delta$. We evaluate their impact using iTransformer model.
(a) Robustness of parameter $\lambda$ on ETTm1.
(b) Effectiveness of the $\delta$ on ETTh1 and ETTm1.
Figure 4: The impact of the hyperparameter on ETTh1 and ETTm1 based iTransformer. More results on parameter $\lambda$ are shown in Appendix F.
PS loss weight $\bm{\lambda}$. The weight $\lambda$ determines the relative importance of PS loss in combination with the MSE. We evaluate a range of $\lambda\in\{0.1,0.3,0.5,0.7,1.0,3.0,5.0,10.0\}$, as depicted in Figure 4(a). The results demonstrate that performance improves as $\lambda$ increases, peaking at $\lambda=3.0$. Notably, the performance remains stable across a broad range of $\lambda$, indicating that PS loss is robust and can be easily applied to various datasets and models.
Patch length threshold $\bm{\delta}$. The patch length threshold $\delta$ limits the maximum patch length to maintain the granularity of comparison. We evaluate $\delta\in\{24,36,48,60\}$ and present the results in Figure 4(b). For ETTh1, where the Fourier-based patch length is smaller than $\delta$, increasing $\delta$ does not affect performance. However, for ETTm1, where the Fourier-based patch length exceeds $\delta$, large $\delta$ results in diminished performance, indicating that overly large patches blur fine-grained structural patterns, and proper patch length is crucial for structural alignment.
5 Conclusion
In this study, we addressed the limitations of traditional point-wise loss functions by introducing the Patch-wise Structural (PS) loss, a novel method that leverages local statistical properties for more precise structural alignment. By integrating patch-level analysis and dynamic weighting strategies, PS loss improves forecasting accuracy and strengthens generalization across multivariate time series models. This innovation establishes a new benchmark for accurately modeling complex time series data. In future work, we will focus on developing a new loss function and model that accounts for both local and global structural dynamics within the prediction.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Datasets
Our study leverages a diverse set of forecasting datasets to evaluate the effectiveness of our loss function across various domains.
  • ETT (Electricity Transformer Temperature): The ETT (Zhou et al., 2021) dataset contains seven features of electricity transformer data collected from two separate counties between July 2016 and July 2018. These datasets vary in granularity, with “h” indicating hourly data and “m” indicating 15-minute intervals. The suffixes “1” and “2” refer to two different regions from which the datasets originated.
  • Weather: The Weather (Zhou et al., 2021) dataset includes 21 meteorological factors recorded every 10 minutes in 2020 at the Max Planck Biogeochemistry Institute’s Weather Station.
  • ECL (Electricity Consuming Load): The ECL (Li et al., 2019) dataset contains hourly electricity consumption data of 321 clients from 2012 to 2014.
  • Exchange: The Exchange (Lai et al., 2018) dataset consists of daily exchange rates from 1990 to 2016 for eight countries: Australia, the United Kingdom, Canada, Switzerland, China, Japan, New Zealand, and Singapore.
We follow the same data processing and train-validation-test split protocol as previous works (Zhou et al., 2021; Wu et al., 2023a; Liu et al., 2024a). The ETT dataset is divided into training, validation, and testing sets in a 12/4/4-month ratio, while other datasets are split in a 7:1:2 ratio. The details of the datasets are provided in Table 6.
Table 6: Descriptions of the forecasting datasets. “Channel” shows the variate number of each dataset. “Dataset Size” indicates the size of the (Train, Validation, and Test) split respectively. “Frequency” denotes the sampling interval of time points.
Dataset Channel Length Datset Size Frequency Information
ETTh1 7 17420 (8545, 2881, 2881) 1 hour Electricity
ETTh2 7 17420 (8545, 2881, 2881) 1 hour Electricity
ETTm1 7 69680 (34465, 11521, 11521) 15 min Electricity
ETTm2 7 69680 (34465, 11521, 11521) 15 min Electricity
Weather 21 52696 (36792, 5271, 10540) 10 min Weather
ECL 321 26304 (18317, 2633, 5261) 1 hour Electricity
Exchange 8 7588 (5120, 665, 1422) 1 day Exchange Rates
Appendix B Backbone Models
We select seven state-of-the-art (SOTA) models from different architectures as backbones. These include Transformer-based models: iTransformer (Liu et al., 2024a) and PatchTST (Nie et al., 2023); MLP-based models: DLinear (Zeng et al., 2023) and TimeMixer (Wang et al., 2024b); CNN-based model: TimesNet (Wu et al., 2023a); LLM-based models: OFA (Zhou et al., 2023) and AutoTimes (Liu et al., 2024b). We use the official implementations of each backbone from their respective repositories and follow their optimal hyperparameter settings for our experiments.
Appendix C Related Work
Patching Mechanism in Time Series Forecasting. In recent years, patch-based approaches have gained significant traction in time series forecasting, offering substantial improvements in prediction performance (Zhang & Yan, 2023; Chen et al., 2024b; Tang & Zhang, 2024). By segmenting the input time series into patches, these methods enhance locality and capture comprehensive semantic information that point-level techniques often overlook (Nie et al., 2023).
For instance, PatchTST (Nie et al., 2023) employs series patching combined with a channel-independent (CI) modeling strategy to capture long-term temporal dependencies. Crossformer (Zhang & Yan, 2023) partitions time series into patches and utilizes self-attention mechanisms to model dependencies across both temporal and variable dimensions. TimesNet (Wu et al., 2023a) and PDF (Dai et al., 2024) apply Fast Fourier Transform (FFT) to determine the dominant frequency of the input series, segmenting the data accordingly to ensure patches capture clear periodic patterns. HDMixer (Huang et al., 2024b) adaptively expands patch lengths to enrich boundary information and mitigate semantic incoherence within the series.
While these methods employ patching to extract localized semantic information from input time series, we take a novel approach by introducing patching to the prediction series. This enables a localized comparison between prediction and ground truth, providing a more fine-grained assessment of structural similarity.
Appendix D Running Cost
To evaluate the computational cost introduced by PS loss, we measure the running time per iteration (ms/iter) of backbone models trained with MSE loss and PS loss across different forecasting lengths on the ETTh1 dataset. As shown in Table 7, although PS loss results in a modest increase in computation time, it consistently enhances forecasting accuracy across all models. Notably, DLinear achieves the highest average MSE improvement of 6.1%, highlighting the efficiency-accuracy trade-off introduced by PS loss.
Table 7: Running cost of PS loss. The table reports the running time per iteration (ms/iter) for backbone models using MSE loss and PS loss on the ETTh1 dataset. “Time Inc.” represents the additional running time introduced by PS loss, while “MSE Dec.” indicates the forecasting performance improvement achieved with PS loss.
Models iTransformer Dlinear TimesNet
Running Cost MSE Loss PS Loss Time Inc. MSE Dec. MSE Loss PS Loss Time Inc. MSE Dec. MSE Loss PS Loss Time Inc. MSE Dec.
(ms/iter) (ms/iter) (ms/iter) (%) (ms/iter) (ms/iter) (ms/iter) (%) (ms/iter) (ms/iter) (ms/iter) (%)
ETTh1 96 22.9 30.4 7.6 2.1% 12.5 19.3 6.9 4.4% 65.2 70.9 5.7 0.6%
192 24.7 31.3 6.6 2.9% 12.0 19.1 7.1 9.2% 66.5 74.9 8.3 4.5%
336 26.5 32.1 5.6 3.4% 12.2 19.4 7.2 2.7% 84.3 92.7 8.4 7.5%
720 27.7 31.7 4.0 5.7% 14.2 21.3 7.1 8.2% 107.5 118.4 10.9 5.4%
Avg 25.4 31.3 5.9 3.5% 12.7 19.7 7.1 6.1% 80.8 89.2 8.3 4.5%
Appendix E Impact of PS Loss on Generalization
To examine how PS loss affects training dynamics and generalization, we visualize the MSE on both the training and test datasets across all epochs, using MSE loss and PS loss as objective functions, as shown in Figure 6. The experiments were conducted on the ETTh1, ETTh2, ETTm1, and ETTm2 datasets, with iTransformer as the backbone model.
(a) Results on ETTh1 dataset.
(b) Results on ETTh2 dataset.
(c) Results on ETTm1 dataset.
(d) Results on ETTm2 dataset.
Figure 6: Training and testing MSE loss curves across all training epochs for the iTransformer model trained with MSE loss and PS loss on the ETT datasets (left: Training errors, right: Testing errors). The model trained with PS loss shows higher training errors but achieves lower testing errors, highlighting its effectiveness in enhancing generalization and mitigating overfitting.
From Figure  6, we observe a consistent trend across both datasets. During training, models optimized solely with MSE loss achieve lower training error per epoch compared to those optimized with PS loss. However, on testing data, models trained with MSE loss exhibit higher testing error per epoch compared to those trained with PS loss. These observations suggest that models trained with MSE loss focus solely on the point-wise error, resulting in smaller training losses, but fail to generalize effectively to the testing data. In contrast, PS loss mitigates overfitting by encouraging the model to focus on local structural similarity in addition to point-wise accuracy, highlighting its effectiveness in improving generalization and forecasting accuracy.
Appendix F More Results on Hyperparameter Sensitivity
We investigate the impact of hyperparameter PS loss weight $\lambda$ on ETTm2 and Weather datasets using iTransformer and DLinear as the backbone models. We set $\lambda\in\{0.1,0.3,0.5,0.7,1.0,3.0,5.0,10.0\}$ for the experiment.
(a) iTransformer as backbone. (left: ETTm2, right: Weather)
(b) DLinear as backbone. (left: ETTm2, right: Weather)
Figure 8: Influence of PS loss weight $\lambda$ on iTransformer and DLinear backbone. We plot the change in MSE with respect to $\lambda$ on the ETTm2 and Weather datasets, using iTransformer and DLinear as the backbone model. The performance is robust across a wide range of $\lambda$ values.
The results show that as $\lambda$ increases, the performance of the model improves, with $\lambda=5.0$ achieving the relatively best results. Moreover, the performance remains robust across a wide range of $\lambda$ values tested, which simplifies PS loss application to different datasets and models.
Appendix G Forecasting Visualization
To evaluate the quality of predictions using MSE loss and our proposed PS loss, we visualize samples from the ETTm1, ETTm2, and Weather datasets with iTransformer as the backbone model. Predictions generated with PS loss exhibit enhanced alignment with the ground truth, accurately capturing directionality, fluctuation dynamics, and average levels while preserving localized patterns and structural details.
(a) Results on ETTm1 dataset.
(b) Results on ETTm2 dataset.
(c) Results on Weather dataset.
Figure 10: Forecasting visualization comparing PS loss and MSE loss as objective functions.
Appendix H More Results on Ablation Study
We conduct additional ablation studies on the DLinear and iTransformer backbones using the ETTh1, ETTh2, ETTm1, and ETTm2 datasets to further demonstrate the necessity of each component, as shown in Table 8 and Table 9.
Table 8: Ablation study of the components of PS loss on the ETT datasets using DLinear as a backbone.
Method Dlinear + PS w/o $\mathcal{L}_{Corr}$ w/o $\mathcal{L}_{Var}$ w/o $\mathcal{L}_{Mean}$ w/o Patching w/o Weighting
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.367 0.389 0.381 0.402 0.380 0.399 0.372 0.395 0.380 0.401 0.384 0.404
192 0.402 0.411 0.439 0.446 0.440 0.444 0.404 0.414 0.436 0.441 0.403 0.411
336 0.435 0.435 0.443 0.445 0.439 0.486 0.440 0.443 0.442 0.442 0.439 0.440
720 0.463 0.484 0.510 0.519 0.509 0.517 0.474 0.493 0.522 0.528 0.528 0.532
Avg 0.417 0.430 0.443 0.453 0.442 0.462 0.423 0.436 0.445 0.453 0.439 0.447
ETTh2 96 0.280 0.341 0.289 0.351 0.281 0.342 0.287 0.351 0.280 0.341 0.283 0.345
192 0.358 0.396 0.381 0.415 0.359 0.396 0.397 0.428 0.386 0.399 0.371 0.404
336 0.435 0.450 0.455 0.466 0.457 0.462 0.477 0.477 0.438 0.450 0.439 0.452
720 0.597 0.540 0.720 0.599 0.673 0.574 0.747 0.612 0.663 0.583 0.679 0.580
Avg 0.417 0.432 0.461 0.458 0.443 0.443 0.477 0.467 0.442 0.443 0.443 0.445
ETTm1 96 0.296 0.339 0.300 0.343 0.296 0.338 0.301 0.345 0.304 0.347 0.297 0.340
192 0.333 0.362 0.334 0.364 0.332 0.359 0.335 0.366 0.341 0.373 0.332 0.361
336 0.365 0.380 0.370 0.387 0.368 0.381 0.369 0.386 0.371 0.387 0.367 0.381
720 0.419 0.413 0.424 0.420 0.422 0.415 0.423 0.420 0.423 0.419 0.422 0.415
Avg 0.353 0.374 0.357 0.378 0.355 0.373 0.357 0.379 0.360 0.381 0.355 0.374
ETTm2 96 0.163 0.251 0.170 0.265 0.163 0.252 0.171 0.268 0.164 0.254 0.164 0.255
192 0.222 0.296 0.234 0.308 0.222 0.296 0.238 0.320 0.223 0.299 0.225 0.301
336 0.277 0.332 0.282 0.342 0.283 0.342 0.295 0.353 0.286 0.347 0.289 0.350
720 0.377 0.397 0.414 0.432 0.398 0.418 0.451 0.453 0.393 0.415 0.410 0.426
Avg 0.260 0.319 0.275 0.337 0.267 0.327 0.289 0.348 0.266 0.329 0.272 0.333
[…]
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.379 0.396 0.381 0.399 0.385 0.400 0.382 0.397 0.383 0.400 0.380 0.396
192 0.428 0.424 0.441 0.433 0.436 0.429 0.431 0.426 0.435 0.429 0.430 0.425
336 0.474 0.453 0.486 0.458 0.499 0.466 0.471 0.446 0.483 0.455 0.485 0.458
720 0.480 0.478 0.515 0.496 0.521 0.498 0.477 0.471 0.511 0.492 0.532 0.505
Avg 0.440 0.438 0.456 0.447 0.460 0.448 0.440 0.435 0.453 0.444 0.457 0.446
ETTh2 96 0.289 0.340 0.294 0.344 0.296 0.343 0.292 0.343 0.290 0.340 0.292 0.341
192 0.373 0.391 0.374 0.393 0.375 0.392 0.379 0.395 0.376 0.393 0.374 0.392
336 0.416 0.428 0.423 0.431 0.421 0.429 0.424 0.431 0.419 0.428 0.419 0.428
720 0.421 0.441 0.439 0.453 0.425 0.442 0.426 0.443 0.423 0.441 0.423 0.441
Avg 0.375 0.400 0.383 0.405 0.379 0.402 0.380 0.403 0.377 0.401 0.377 0.401
ETTm1 96 0.326 0.360 0.334 0.365 0.328 0.360 0.350 0.375 0.343 0.374 0.328 0.361
192 0.374 0.384 0.382 0.387 0.374 0.381 0.384 0.390 0.385 0.394 0.375 0.383
336 0.410 0.406 0.419 0.412 0.412 0.406 0.417 0.412 0.420 0.416 0.411 0.407
720 0.472 0.440 0.482 0.446 0.479 0.443 0.483 0.449 0.488 0.455 0.477 0.444
Avg 0.396 0.397 0.404 0.402 0.398 0.397 0.408 0.407 0.409 0.410 0.398 0.399
ETTm2 96 0.175 0.254 0.179 0.259 0.177 0.256 0.183 0.266 0.182 0.265 0.177 0.258
192 0.242 0.299 0.248 0.306 0.244 0.300 0.247 0.306 0.248 0.306 0.245 0.302
336 0.304 0.338 0.309 0.343 0.308 0.340 0.308 0.344 0.310 0.344 0.309 0.343
720 0.401 0.394 0.409 0.400 0.405 0.397 0.410 0.401 0.410 0.400 0.410 0.399
Avg 0.281 0.321 0.286 0.327 0.283 0.323 0.287 0.329 0.287 0.329 0.285 0.325
  • Loss components. Omitting any one of the three loss components results in a decline in performance, underscoring the unique and complementary roles each loss plays in capturing localized structural similarities.
  • Adaptive Patching. Comparing time series globally leads to performance degradation, especially for longer forecasting horizons. This highlights the importance of point-wise comparison in capturing nuanced local variations in statistical properties.
  • Dynamic weighting. Replacing the GDW strategy with fixed weighting leads to performance degradation, highlighting the importance of GDW in ensuring the balanced utilization of each loss term for effective model training.